OpenStudy (anonymous):

decimal as percant 0.003 decimal as percant 0.003 @Mathematics

OpenStudy (saifoo.khan):

0.3%

OpenStudy (anonymous):

how

OpenStudy (saifoo.khan):

% = 1/100 = 0.003/100 = 0.3%